The line passing through (7,y) and (1,0) is parallel to the line joining (4,4) and (-2,1) Find y
Answer by richwmiller(17219)

first you need to find the slope of the line joining (4,4) and (-2,1)
the line through (7,y) and (1,0) will have the same slope.
(y-y)/(x-x)=m
1-4/-2-4=m
-3/-6=m
1/2=m
(0-y)/(1-7)=1/2
-y/-6=1/2
-y=-6/2
y=6/2
y=3